OpenStudy (amonoconnor):

I've been doing pretty good with these Definite Integrals, but this one has me stumped. Can someone help me understand how to work through this? The problem: g(x) = (integral)[(u^2 - 1)/(u^2 + 1)]du, from 2x to 3x. Any and all help is greatly appreciated!

OpenStudy (amonoconnor):

The problem: \[g(x) = \int\limits_{2x}^{3x}(\frac{u^2 - 1}{u^2 + 1})du\]

jimthompson5910 (jim_thompson5910):

Does this hint help? w = u^2 + 1 w-2 = u^2 + 1-2 w-2 = u^2 - 2

jimthompson5910 (jim_thompson5910):

hmm now that I think of it, if we derived both sides of the equation with respect to u, then dw/du = 2u dw = 2u*du but there is no u*du term in the integrand

OpenStudy (amonoconnor):

I'm not really sure where to go with your hint @jim_thompson5910 :/

OpenStudy (superdavesuper):

i cheated n checked out the integrand in Wolfram ;) u can use the substitution u=tan(theta)

jimthompson5910 (jim_thompson5910):

that trick will work

jimthompson5910 (jim_thompson5910):

u = tan(theta) du/dtheta = sec^2(theta) du = sec^2(theta)*dtheta u^2 + 1 = tan^2(theta) + 1 = sec^2(theta) so we'll have a pair of "sec^2(theta)" terms cancel

OpenStudy (amonoconnor):

Wouldn't it be inverse tan(x)? As in, \[\frac{d}{dx}\tan^{-1} x = \frac{1}{1 + x^2}\]

OpenStudy (superdavesuper):

correct @jim_thompson5910 and @amonoconnor

OpenStudy (amonoconnor):

What about those boundaries... Do I need to split it up into two integrals, with 2x as the upper for one and 3x as the upper for another? Can 3x=b and 2x=a and I just use it as is?

OpenStudy (superdavesuper):

well u=tan(theta) so finish the integrand in theta first; then subsitute u back in; last evaluate u=2x and 3x

OpenStudy (amonoconnor):

Can you walk me through this? I'm going to try on my end...

OpenStudy (amonoconnor):

Do I do it like a chain rule? Plugging it in, then multiply the "new" integrand by d/dx[tan(theta)]

OpenStudy (superdavesuper):

promise u wont click till u try to solve the intregrand 1st: http://www.wolframalpha.com/input/?i=integrate+%28u%5E2++-+1%29%2F%28u%5E2+%2B+1%29du

OpenStudy (amonoconnor):

Deal.

OpenStudy (superdavesuper):

follow the steps in the last post by @jim_thompson5910 and substitute u and du into the integrand

OpenStudy (amonoconnor):

Here's what I'm understanding... tell me if I'm wrong! \[g(x) = \int\limits_{2x}^{3x}\frac{u^2-1}{u^2+1}*du\] \[= \int\limits_{2x}^{3x}\frac{u^2+1}{1}*\frac{-1}{u^2+1}*du\] \[f(x) = \frac{u^2+1}{1}*\frac{-1}{u^2+1}\] \[F(x) = \frac{1}{\tan(x)}*(-\tan(x))\]

jimthompson5910 (jim_thompson5910):

you can't jump from \[\Large g(x) = \int\limits_{2x}^{3x}\frac{u^2-1}{u^2+1}*du\] to \[\Large \int\limits_{2x}^{3x}\frac{u^2+1}{1}*\frac{-1}{u^2+1}*du\]

jimthompson5910 (jim_thompson5910):

\[\Large \frac{u^2-1}{u^2+1} \ne \frac{u^2+1}{1}*\frac{-1}{u^2+1}\]

OpenStudy (amonoconnor):

Ugg. Thank you, I'll keep working...

OpenStudy (superdavesuper):

it should go like: u=tan(th), du=sec^2(th)d(th) so the original integrand becomes: (tan^2(th)-1)/(tan^2(th)+1) sec^2(th)d(th) sorry im super lazy in typing things out in latex!

OpenStudy (amonoconnor):

Alright... tell me if this is right:) \[g(x) = \int\limits_{2x}^{3x}\tan^2(x)-1\] And from there, I evaluate by: \[F(b) - F(a)\] ???

OpenStudy (superdavesuper):

forget x, g(x) and focus on u and theta first....

OpenStudy (amonoconnor):

plugging in 2x and 3x?

OpenStudy (superdavesuper):

no...forget the x part completely. for now - solve the integrand in u by substitution w/ theta

jimthompson5910 (jim_thompson5910):

\[\Large u = \tan(\theta)\] \[\Large \frac{du}{d\theta} = \sec^2(\theta)\] \[\Large du = \sec^2(\theta)*d\theta\] ------------------------------------------------------------------------- \[\Large \int\left(\frac{u^2 - 1}{u^2 + 1}\right)du\] \[\Large \int\left(\frac{(\tan(\theta))^2 - 1}{(\tan(\theta))^2 + 1}\right)\sec^2(\theta)*d\theta\] \[\Large \int\left(\frac{\tan^2(\theta) - 1}{\tan^2(\theta) + 1}\right)\sec^2(\theta)*d\theta\] \[\Large \int\left(\frac{\tan^2(\theta) - 1}{\sec^2(\theta)}\right)\sec^2(\theta)*d\theta\] \[\Large \int\left(\frac{\tan^2(\theta) - 1}{\cancel{\sec^2(\theta)}}\right)\cancel{\sec^2(\theta)}*d\theta\] \[\Large \int\left(\tan^2(\theta) - 1\right)d\theta\]

OpenStudy (amonoconnor):

Gotcha... okay, what's after that?

OpenStudy (superdavesuper):

ty @jim_thompson5910

OpenStudy (superdavesuper):

@amonoconnor continue w/ the integration!

jimthompson5910 (jim_thompson5910):

use the identity tan^2(x) + 1 = sec^2(x) tan^2(x) + 1-1 = sec^2(x)-1 tan^2(x) = sec^2(x) - 1

OpenStudy (amonoconnor):

Why do I need to go back to sec^2(x) stuff, if I have tan^2(theta) - 1 as the only term under the integral?

jimthompson5910 (jim_thompson5910):

because tan^2 is a pain to integrate compared to sec^2

jimthompson5910 (jim_thompson5910):

y = tan(x) dy/dx = sec^2(x) you have to think backwards a bit

OpenStudy (amonoconnor):

Gotcha.... so, from there, I need to find F(x), THEN plug in boundaries?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (amonoconnor):

Alright, I'm doing better... let me work this step!

OpenStudy (amonoconnor):

f(x) = sec^2(theta) - 2 F(x) = tan(theta) - 2(theta) ???

OpenStudy (amonoconnor):

Just checking about that 2.

jimthompson5910 (jim_thompson5910):

I agree with the `tan(theta) - 2(theta)` part

jimthompson5910 (jim_thompson5910):

now go back to `u = tan(theta)` and isolate theta to get `theta = arctan(u)`

OpenStudy (amonoconnor):

Wait... what? \[F(x) = \tan(\tan^{-1}(u)) - 2(\tan^{-1}(u))\] ???

OpenStudy (amonoconnor):

I thought I had gotten the antiderivative already? :/

OpenStudy (amonoconnor):

What is tan of arctan(u)?

jimthompson5910 (jim_thompson5910):

the limits of integration (2x and 3x) are in terms of u ie, u = 2x is the lowest u goes u = 3x is the highest u goes which is why I'm resubstituting theta for an expression in terms of u

jimthompson5910 (jim_thompson5910):

tan(arctan(u)) = u

jimthompson5910 (jim_thompson5910):

they are inverses of each other, so in a sense, they cancel out

OpenStudy (amonoconnor):

I'm so sorry for asking a lot of questions, but I'm just struggling on this one... So.. 2(theta) = 2(tan^-1(u)) = 2(tan^-1(3x)) = ?

OpenStudy (amonoconnor):

I get how tan(tan^-1(u)) = u.

OpenStudy (superdavesuper):

@amonoconnor alright - click on the wolfram link in my earlier post. it gives u the integrand in terms of u. no more theta to create any more confusion ;) then substitute u=2x and u=3x to get the ans for g(x). good luck!

jimthompson5910 (jim_thompson5910):

you should find that \[\Large \int\left(\frac{u^2 - 1}{u^2 + 1}\right)du = u-2\arctan(u) + C\]

jimthompson5910 (jim_thompson5910):

now plug in the limits of integration \[\Large \int_{\color{red}{2x}}^{\color{blue}{3x}}\left(\frac{u^2 - 1}{u^2 + 1}\right)du \\ \Large \left. = u-2\arctan(u) + C\right]_{\color{red}{2x}}^{\color{blue}{3x}}\\ \Large = \left[\color{blue}{3x}-2\arctan(\color{blue}{3x}) + C\right]-\left[\color{red}{2x}-2\arctan(\color{red}{2x}) + C\right]\\ \Large =??\]

jimthompson5910 (jim_thompson5910):

Or you can think of it as \[\Large h(u) = u-2\arctan(u) + C\] and then compute `h(3x) - h(2x)` to finish up

OpenStudy (amonoconnor):

x - 2(arctan(3x) - arctan(2x)) ?

jimthompson5910 (jim_thompson5910):

that or \[\Large g(x) = x - 2\arctan(3x) + 2\arctan(2x)\]

OpenStudy (amonoconnor):

Alright, cool.. I'm tired, and will likely look back at this one in the morning and have more enlightenment.. Thank you for all of your patient help:)

OpenStudy (superdavesuper):

special thanx to @jim_thompson5910 - i would have given u 5 medals if i can :)

OpenStudy (amonoconnor):

Ditto! You are both Rockstars! ;)